# Real-time 3D-aware Portrait Editing from a Single Image

### v2

@@ 5 Discussion
Limitations. Despite achieving state-of-the-art performance in terms of quality and efficiency, our method exhibits inconsistencies in details for novel-view rendering because the EG3D framework relies on a super-resolution module. Additionally, when our method is applied to video editing, it presents flickering artifacts since our model is originally designed for image-based editing.
@@
Conclusion. In this work, we introduce 3DPE for real-time 3D-aware portrait editing. By distilling the powerful knowledge of diffusion models and 3D GANs into a lightweight module, our model significantly reduces editing time while ensuring quality. Benefiting from the distillation process, 3DPE could handle multiple styles in a single model and is able to perform both image-guided and text-guided editing.
Appendix
Appendix 0.A Overview
This appendix is organized as follows. Appendix 0.B describes the dataset and demonstrates data samples of the experiments. Appendix 0.D provides additional 3D-aware editing results including comparisons with prior baselines and customized editing results. Appendix 0.E further analyzes the representation of Live3D with experimental results. The final section describes the interactive system mentioned in the main text.
Appendix 0.B Dataset
Here we provide samples of the training set for investigation in Fig. 8. As described in the main text, we adopt the text-guided 2D editing model [7] as the data engine. During training, the input will be sent to the model accompanied by a randomly sampled prompt. The pseudo-labels generated by [7] are adopted for further supervision, as described in the main text.
Figure 8: The visualization of training samples.
Appendix 0.C Interactive System
Figure 9: The interactive system. By clicking the ‘submit’ button, the system will automatically apply our method on the given image and prompt, and output the edited images as well as novel-view edited results.
We design an interactive interface for users to experience efficient editing, which is shown in Fig. 9. Users are required to provide a prompt to indicate the desired style and an input image that is to be stylized. By clicking the ‘submit’ button, the system will automatically apply our method on the given image and prompt, and output the edited images as well as novel-view edited results.
Appendix 0.D Additional 3D-aware Editing Results
We provided additional qualitative comparisons on 3D-aware editing. The baselines include PTI [54] + StyleCLIP [49], SPI [79] + StyleGAN-NADA [21], and InstructPix2Pix (IP2P) [7] + Live3D [68]. As in Fig. 10, PTI + StyleCLIP generates low-fidelity textures and geometry. SPI + StyleGAN-NADA performs poorly in following the prompts. As demonstrated by the former three cases, IP2P + Live3D finds it hard to perceive the image prompts and fails in editing guided by images, especially for the styles hard to be described with texts. For text prompts, our method outperforms IP2P + Live3D in identity preservation and multi-view consistency (note the eye gaze in the fourth case). Generally, our method outperforms the baselines in both quality and efficiency.
Additional qualitative results are demonstrated in Fig. 11 and we further provide customized editing results in Fig. 12, where users only need to provide 2D image pairs to customize their own style. Note that in the experiments, all of the input images and reference images (image prompts) have not been shown to the model during training. And the experimental results reveal that our method could well adapt to the user-defined customized styles.
Figure 10: Additional qualitative comparisons. We compare the results of several baselines with image prompts and text prompts. In each case, we include the edited portraits as well as their novel view renderings. Our method outperforms in both quality and efficiency.
Figure 11: Additional qualitative results.
Figure 12: Additional customized editing results. The prompts in this experiment have not been shown to the model in the prior training.
Appendix 0.E Additional Validation of Live3D Features
Recall that in the main text, we analyze the feature representation of Live3D, which serves as the motivation for our training strategy. In this section, we provide additional results. The experimental setting of Fig. 13 is the same as the main text. In each example, we demonstrate 3 lines of 3D reconstruction results for various versions of Live3D - the original full model, the model without $E_{high}$, and the model without $E_{low}$, where “w/o” means we randomize the weights of the module. We can conclude from the results that the features of $E_{low}$ and $E_{high}$ separately model low-frequency and high-frequency information.
To further validate this property, we conduct experiments on the video sequence. We first disable $E_{low}$ and send three frames in the video to Live3D for reconstruction. As in Fig. 14, the eyes of the input portrait are blue and the reflections seem purple. And the reconstructed images only preserve the texture and the texture color is highly similar to the one of the reflections on the glasses and the eyes. Note that when the input portrait closes his eyes, the color of the reconstructed textures becomes purple, which further validates that features of $E_{high}$ tend to model high-frequency information such as color.
Figure 13: Additional experimental results on Live3D feature representation. The features of $E_{low}$ and $E_{high}$ separately tend to model low-frequency and high-frequency information.
Figure 14: Experimental results for the video sequence on Live3D feature representation. When the input portrait closes his eyes, the color of the reconstructed textures becomes purple, which further validates that features of $E_{high}$ tend to model high-frequency information such as color.

### v3

@@ 5 Discussion
Limitations. Despite achieving state-of-the-art performance in quality and efficiency, our method exhibits inconsistencies in details for novel-view rendering because the EG3D framework relies on a super-resolution module. Additionally, when our method is applied to video editing, it presents flickering artifacts since our model is originally designed for image-based editing.
@@
Conclusion. In this work, we introduce 3DPE for real-time 3D-aware portrait editing. By distilling the powerful knowledge of diffusion models and 3D GANs into a lightweight module, our model significantly reduces editing time while ensuring quality. Benefiting from the distillation process, 3DPE could handle multiple styles in a single model and perform both image-guided and text-guided editing. The advantages of our method empower us to generate photorealistic 3D portraits, a capability crucial for the visual effects industry, AR/VR systems, and teleconferencing, among other applications, while we firmly oppose any misuse that could compromise privacy and security.
Acknowledgements
This project was supported by the National Key R&D Program of China under grant number 2022ZD0161501.